Please help me solve this equation:
The product of two consecutive odd integers is equal to 30 more than the first. Find the integers.
Thank you!

Answer by TimothyLamb(4379)

---
y = x + 2
xy = x + 30
---
xy = x + 30
x(x + 2) = x + 30
xx + 2x = x + 30
xx + x - 30 = 0
---
the above quadratic equation is in standard form, with a=1, b=1, and c=-30
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 1 -30
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 5
x = -6
---
the negative root doesn't fit the problem statement, so use the positive root:
---
solution:
x = 5
y = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php